OpenStudy (anonymous):

Anyone in Keystone National High School Algebra 2? Really need help!

OpenStudy (anonymous):

I am, what are you doing now?

OpenStudy (anonymous):

Consistent and inconsistent quiz, but pretty much the part two of the Algebra course.

OpenStudy (anonymous):

ok so I'll take the first 2 for an example. 1. Is the following system consistent or inconsistent? 8x + y + 4 = 0 y = -8x - 4 that's consistent because it has all the same numbers pretty much. 2. Is the following system consistent or inconsistent? 7x + 6 = -2y -14x -4y + 2 = 0 that would be inconsistent because it doesn't stay the same and isn't increased by the same factor, does that make sense?

OpenStudy (anonymous):

A little can you explain the second one that is inconsistent a little more?

OpenStudy (anonymous):

yeah, so the 7x is increasing at a rate of -2, making it -14x, but the -2y is increasing by 2, bot -2. Also, you would have all of the normal numbers like the 6 would all be moved to one side, plus it isn't increased or decreased by -2 or 2, so everything is inconsistently changed. Make more sense?

OpenStudy (anonymous):

So, here's another example.. y = 2x + 5 -2x + y = -2 so this will be inconsistent because you're not keeping the 5 a five when y is staying the same and the 2x becomes -2x because you are subtracting it on both sides. the -2 on the right side of the equation would still be 5.

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

no problem :) let me know if you have other questions in the unit

OpenStudy (anonymous):

Okay, just curious are you finished with the course?

OpenStudy (anonymous):

I'm in last unit, taking the final tomorrow.